Spring balance reads 10kg in a bucket of water is suspended from it what will be the reading of the balance when an iron piece of mass 7.2 kg suspended by a string is immersed  within half its volume inside the water in the bucket relative density of the iron is 7.2

Grade:

1 Answers

Arun
5 years ago
Mass of iron piece = 7.2 kg
Density of iron = Mass × Volume 
Volume of iron piece = 7.2/7.2×10 = 10^-3 kg^3
Volume of iron immersed = 10^-3/2 = 
Mass of water displaced by iron piece = 10^-3/2 × 10^3 = 0.5 kg
Now, reading in spring balance = 10+0.5 = 10.5kg